# A Simulation Based Inference Approach to the Dynamics of the MW–LMC System - Validation

### v1

@@ 4.3 On a cosmological MW–LMC analogue
Figure 4: Validation on a cosmological MW–LMC analogue: The joint, and individual, posterior distributions for MW mass, LMC mass and dynamical friction strength when there is knowledge of the LMC present day position, velocity and the reflex motion parameters (purple). We also show the parameter prior distributions in grey. For the 1D posterior panels we show the $16^{\rm th}$ and $84^{\rm th}$ percentiles as shaded regions. The yellow squares/lines represent the true values of m12b’s MW and LMC analogue from the FIRE latte project (Wetzel et al., 2023). The all-sky projection displays the average solar corrected radial velocity, $v_{\rm GSR}$, for all stellar halo particles beyond $50\,\rm kpc$. The net signal across the entire sky is subtracted. The pink cross denotes the reflex motion apex for this simulation. The past orbit of the LMC in this potential is shown as the black dashed line. Orbits computed using posterior samples for the masses and dynamical friction strength, for fixed LMC present-day conditions, are shown as the grey shaded region.
@@
To demonstrate this, we use a MW analogue from the Latte suite within the FIRE (Wetzel et al., 2023) project [^3] Feedback In Realistic Environments webpage. These simulations are zoom-in cosmological-hydrodynamical simulations run using the FIRE-2 model (see, Hopkins et al., 2018, for a detailed description of the full physics model). We select the m12b simulation from the Latte suite as this is the closest analogue to the real MW-–LMC system (see, Garavito-Camargo et al., 2024, table 1 for a summary of the Latte MW–LMC analogue properties). This simulation produces a present-day MW with a stellar mass, gas mass and the dark matter mass / density profile that are similar to the real MW (Hopkins et al., 2018; Garrison-Kimmel et al., 2018; Sanderson et al., 2020). The present-day mass of the MW analogue is $M_{200,\rm MW}=9.0\times 10^{11}\,\rm M_{\odot}$. Meanwhile, the LMC analogue in m12b has an infall mass of $M_{\rm LMC}=21\times 10^{10}\,\rm M_{\odot}$ such that its mass ratio to the MW host is $\sim 1:4$.
@@
The Latte suite of simulations are carried out in an arbitrary simulation box frame. However, in order to take advantage of our rigid MW–LMC simulation set described in Sec. 2, we must rotate this cosmological analogue to have the same orientation as the real MW–LMC system. To achieve this, we follow the same procedure as in appendix A of Arora et al. (2024) whereby we first establish a principal axis that aligns the galactic disc to the $x-y$ plane at the epoch of the analogue’s first pericentric passage. Thereafter, an additional rotated axis is defined to align the position of the LMC analogue at the epoch of first pericentric passage with that of the real LMC. To do this, the real pericentre of the LMC is based on the orbits in Garavito-Camargo et al. (2019). Applying these rotational transformations to m12b’s LMC analogue gives a present-day position of $\mathbf{x}_{\mathrm{LMC}}=\{1.8,-16.7,-34.0\}\,\rm kpc$ and a velocity of $\mathbf{v}_{\mathrm{LMC}}=\{-87.7,-313.4,149.4\}\,\rm km\,\rm s^{-1}$. The same rotational transformations are applied to the stars of the m12b MW analogue. In this frame, we can measure the MW reflex parameters using the method outlined in Sec. 2.2. The stars of the m12b MW analogue have been masked for any substructures that could lead to biasing of the returned parameters. In this context, substructure is defined as any stars that are associated with satellite galaxies and the LMC analogue. Again, we ensure that the number of particles approximately matches the average number density of stars beyond $50\,\rm kpc$ in the DESI BHB sample within the DESI sky-coverage footprint (Byström et al., 2024, adopting equ. 11 for the DESI footprint). The resulting reflex motion parameters as measured from this simulation are $v_{\mathrm{travel}}=50.0\,\rm km\,\rm s^{-1},l_{\mathrm{apex}}=-8.97^{\circ},b_{\mathrm{apex}}=-73.4^{\circ}$; see the all-sky projection in Fig. 4 for the direction of the apex on-sky (pink cross).
@@
We look to estimate the MW analogue mass, LMC analogue mass and the dynamical friction strength with knowledge of the LMC present-day position, velocity and the reflex motion parameters. Using our amortised posterior this can be efficiently evaluated for the given data. In Fig. 4 we show the joint, and individual, posterior distributions. Interestingly, our SBI framework is able to recover the true LMC mass from the cosmological MW–LMC analogue within the $16^{\rm th}$ and $84^{\rm th}$ percentiles. However, the mean of the LMC mass posterior distributions is biased towards higher values. Moreover, the dynamical friction strength is well constrained to values that are $\sim 2-3$ times greater than, and intriguingly not consistent with, those expected by the classical Chandrasekhar dynamical friction (i.e., $\log_{10}(\lambda_{\rm DF})=0$, Chandrasekhar, 1943). Similar to the deforming simulations, no specific value for the ‘true’ dynamical friction strength is quoted for this cosmological simulation, hence its absence from Fig. 4. Again we can use the orbit of this LMC analogue to find an approximate value given our treatment of dynamical friction which returns a value consistent with the posterior in Fig. 4. Exploring the ability to use our SBI framework to constrain dynamical friction remains of interest for future studies. Once more, the MW mass posterior is not well constrained away from the original prior distribution, implying that the supplied data points do not break model degeneracies.
@@ 5.1 Coverage probability test
Any posterior resulting from a generative model should be assessed for its accuracy through a variety of diagnostic tests before being used for inference given the actual observed data. The accuracy of an estimated posterior is often performed using some form of a coverage probability test. A coverage test in Bayesian analysis checks whether credible intervals have the expected probabilities (see, Jeffrey et al., 2025, sec. 2.4 for a concise explanation). In a 1-dimensional posterior setting, one can define a particular credible interval to be the narrowest interval containing, for example, $90\%$ of the probability weight. The Bayesian inference procedure takes in some observed data, $D_{\rm obs}$, and determines a posterior distribution, $p(\theta|D_{\rm obs})$, and hence a credible interval for $\theta$. For a coverage test, one uses a test parameter, $\theta_{\rm test}$, selected from the prior, $p(\theta)$, as the input to a simulation that produces the corresponding output data point, $D_{\rm test}$. From this, one can derive a posterior, $p(\theta|D_{\rm test})$, and therefore a credible interval. If the inference process has been correctly implemented then the true test parameter value, $\theta_{\rm test}$, will fall in this credible interval, in this example, $90\%$ of the time. Repeating this test for many sampled $\theta_{\rm test}$, and varying confidence intervals, one can gain confidence that the estimated posterior distributions are indeed correct.
@@
To perform a coverage test on our SBI posteriors, we will follow the ‘Tests of Accuracy with Random Points’ (TARP, Lemos et al., 2023, see their figs. 1&2 for futher intuition). For our application of SBI, this test is relatively straightforward as we have many pre-existing simulations with an amortised inference scheme, i.e., each data evaluation is computationally cheap. TARP coverage probabilities test the accuracy of estimated posteriors by only using samples made from the posterior. This technique is similar to simulation based calibration (Talts et al., 2018) but extends the idea to the full-dimensional posterior space instead of being solely 1-dimensional. We use the implementation of TARP in the sbi python package (Tejero-Cantero et al., 2020) and show our results in Fig. 5.
@@
A valid SBI posterior will have a coverage probability that is similar to any given credibility interval one chooses. Fig. 5 demonstrates that the expected coverage does indeed match the credibility level. This validates our neural posterior estimation as being truly representative of the probability that each of our model parameters has some true value. This can be further quantified in two ways. Firstly, we can compute the area between the ideal TARP curve and our posterior TARP curves for credibility intervals greater than $0.5$; namely the Area To Curve (ATP) value. This number should be close to 00, a value $\gg 0$ indicates an estimated posterior that is too wide, conversely, a value $\ll 0$ indicates that the estimated posterior is too narrow. Secondly, we can calculate the p-value of a Kolmogorov-Smirnov test. The null hypothesis is that the ideal TARP curve and our posterior TARP curve are identical. If this p-value is less than $0.05$, then this null hypothesis is rejected. For our posterior distributions of MW mass, LMC mass, and dynamical friction strength when there is knowledge of the LMC present-day position, velocity, and reflex motion parameters (purple) we report an ATP value of $0.27$ and a Kolmogorov-Smirnov p-value of $1.0$. These values suggest that we are not drastically over-/under-fitting and are not required to reject the posterior.
@@
Figure 5: Tests of Accuracy with Random Points: This diagnostic check acts as a coverage probability test. The probability of true values in the appropriate credible intervals matches the expected coverage probability. This validates our SBI posterior estimation such that the posterior is representative of the probability that each of our model parameters has some true value.
@@ 5.2 Posterior Predictive Check
Although not explicitly a diagnostic test, we carry out a ‘Posterior Predictive Check’ (PPC) to act as a complementary diagnostic test. This test makes use of the fact that if the inference is correct, then any generated data, $D_{\rm pp}$, using simulation parameter values sampled from the posterior, $\theta_{\rm pp}$, should be similar to the observed data, $D_{\rm obs}$ (Lueckmann et al., 2021). A PCC can provide an intuition about any bias introduced in inference e.g., determining whether or not the generated data systematically differ from the observed data used during the inference.
We carry out a PPC for our posterior distributions of the MW mass, LMC mass, and dynamical friction strength when there is knowledge of the LMC present-day position, velocity, and knowledge of the reflex motion parameters. We repeat this for the three cases we consider in this work: the validation on a rigid simulation in Sec. 4.1, the validation on a deforming simulation in Sec. 4.2 and the validation on a cosmological simulation in Sec. 4.3. To do this, we sample values for the MW mass, LMC mass and dynamical friction strength from the posterior evaluated using data in these three scenarios. We then re-run rigid MW–LMC simulations adopting these parameter values and output the reflex motion parameters to compare to the original input reflex motion data. In Fig. 6, we visualise the resulting generated data as coloured $1\,\sigma$ and $2\,\sigma$ contours and the original data (coloured crosses). The 1D panels show the filled $1\,\sigma$ region. Clearly, the input reflex motion data points used to evaluate the posteriors fall well within the $2\,\sigma$ confidence intervals across all fidelity MW–LMC simulations we consider. Furthermore, for the rigid and deforming simulations, the input data is contained within their $1\,\sigma$ confidence interval. This suggests that our return SBI posteriors are not subject to any biases and are truly representative of the input data.
@@ 6.1.1 Improvements to the Milky Way – LMC models
Firstly, we have assumed that the MW is at equilibrium prior to the LMC’s infall. The extent to which the MW has any ‘smooth’ stellar halo is still of interest. The inner halo is made up of many substructures from accreted dwarf galaxies, globular clusters and stellar streams, some of which are well-mixed in phase-space, whereas others retain a degree of coherent structure at present day (e.g., Naidu et al., 2020; Malhan et al., 2022; Dillamore & Sanders, 2025). Similarly, the outer halo contains a plethora or accreted substructures (e.g., the merger with the Sagittarius dSph galaxy, Helmi, 2020; Vasiliev et al., 2021; Hunt & Vasiliev, 2025, and references therein). Accounting for this induced state of disequilibrium in the MW prior to the LMC’s infall is an oversight in our current modelling. One way to incorporate this is to include an overall net compression velocity to the MW stellar halo due to its contraction over time (Chandra et al., 2025; Byström et al., 2024).
@@
Secondly, we assumed that the LMC has recently made its first pericentric passage. However, there is the alternative scenario where the LMC is on its second pericentric passage (Vasiliev, 2024). In this case, the first pericentric passage would cause a degree of disequilibrium in the MW, which we do not account for. However, the first pericentric passage would have occurred around $\sim 5-10\,\rm Gyr$ ago and at a distance $\geq 100\,\rm kpc$. Hence, almost all dynamical disequilibrium induced by any previous passage will have been superseded by its most recent one (Vasiliev, 2024; Sheng et al., 2024).
@@
Other modelling improvements include accounting for any mass changes of either the MW or LMC, triaxiality in the MW stellar haloes and/or distortions. The mass changes are especially important when considering cosmological analogues as both the host and satellite masses can change by a significant fraction throughout their interaction. In turn, this affects the degree of dynamical friction that the infalling satellite is subject to. In our current analytic dynamical friction prescription, this makes it difficult to separate out the effect of a time varying satellite mass and the inferred value of the dynamical friction strength at present-day, $\lambda_{\rm DF}$ (see Sec. 2.1.3 and Koposov et al., 2023). A further future model improvement would be to implement a more careful treatment of time dependent dynamical friction to account for this. Moreover, Chandrasekhar’s formula becomes increasingly ill-posed when considering two body systems of comparable masses (White, 1983; Binney & Tremaine, 2008; Jethwa et al., 2016). Finally, $v_{\rm travel}$ is sensitive to the mass enclosed by the LMC within a given distance rather than the total LMC mass which involves an extrapolation. We aim to reflect this in future work to return posteriors on the LMC mass enclosed.
6.1.2 Connecting with observations
As discussed above, our rigid simulations do not contain any substructures. Hence we are not required to account for the masking of such substructures that are seen in the observed MW stellar halo (e.g., masking the Sagittarius stellar stream in the DESI BHB sample, Byström et al., 2024; Amarante et al., 2024). This makes comparisons to survey data an interesting problem. One could look to investigate the effects of sky coverage and survey completeness on the returned parameter posterior distributions using our SBI methodology; see more in Sec. 6.4.

### v2

@@ 4.3 On a cosmological MW–LMC analogue
Figure 4: Validation on a cosmological MW–LMC analogue: The joint, and individual, posterior distributions for MW mass, LMC mass and dynamical friction strength when there is knowledge of the LMC present day position, velocity and the reflex motion parameters (purple). We also show the parameter prior distributions in grey. For the 1D posterior panels we show the $16^{\rm th}$ and $84^{\rm th}$ percentiles as shaded regions. The yellow squares/lines represent the true values of m12b’s MW and LMC analogue from the FIRE latte project (Wetzel2023). The all-sky projection displays the average solar corrected radial velocity, $v_{\rm GSR}$, for all stellar halo particles beyond $50\,\rm kpc$. The net signal across the entire sky is subtracted. The pink cross denotes the reflex motion apex for this simulation. The past orbit of the LMC in this potential is shown as the black dashed line. Orbits computed using posterior samples for the masses and dynamical friction strength, for fixed LMC present-day conditions, are shown as the grey shaded region.
@@
To demonstrate this, we use a MW analogue from the Latte suite within the FIRE (Wetzel2023) project [^3] Feedback In Realistic Environments webpage. These simulations are zoom-in cosmological-hydrodynamical simulations run using the FIRE-2 model (see, Hopkins2018, for a detailed description of the full physics model). We select the m12b simulation from the Latte suite as this is the closest analogue to the real MW-–LMC system (see, Garavito-Camargo2024, table 1 for a summary of the Latte MW–LMC analogue properties). This simulation produces a present-day MW with a stellar mass, gas mass and the dark matter mass / density profile that are similar to the real MW (Hopkins2018; Garrison-Kimmel2018; Sanderson2020). The present-day mass of the MW analogue is $M_{200,\rm MW}=9.0\times 10^{11}\,\rm M_{\odot}$. Meanwhile, the LMC analogue in m12b has an infall mass of $M_{\rm LMC}=21\times 10^{10}\,\rm M_{\odot}$ such that its mass ratio to the MW host is $\sim 1:4$.
@@
The Latte suite of simulations are carried out in an arbitrary simulation box frame. However, in order to take advantage of our rigid MW–LMC simulation set described in Sec. 2, we must rotate this cosmological analogue to have the same orientation as the real MW–LMC system. To achieve this, we follow the same procedure as in appendix A of Arora2024 whereby we first establish a principal axis that aligns the galactic disc to the xyx-y plane at the epoch of the analogue’s first pericentric passage. Thereafter, an additional rotated axis is defined to align the position of the LMC analogue at the epoch of first pericentric passage with that of the real LMC. To do this, the real pericentre of the LMC is based on the orbits in Garavito-Camargo2019. Applying these rotational transformations to m12b’s LMC analogue gives a present-day position of $\mathbf{x}_{\mathrm{LMC}}=\{1.8,-16.7,-34.0\}\,\rm kpc$ and a velocity of $\mathbf{v}_{\mathrm{LMC}}=\{-87.7,-313.4,149.4\}\,\rm km\,\rm s^{-1}$. The same rotational transformations are applied to the stars of the m12b MW analogue. In this frame, we can measure the MW reflex parameters using the method outlined in Sec. 2.2. The stars of the m12b MW analogue have been masked for any substructures that could lead to biasing of the returned parameters. In this context, substructure is defined as any stars that are associated with satellite galaxies and the LMC analogue. Again, we ensure that the number of particles approximately matches the average number density of stars beyond $50\,\rm kpc$ in the DESI BHB sample within the DESI sky-coverage footprint (Bystrom2025, adopting equ. 11 for the DESI footprint). The resulting reflex motion parameters as measured from this simulation are $v_{\mathrm{travel}}=50.0\,\rm km\,\rm s^{-1},l_{\mathrm{apex}}=-8.97^{\circ},b_{\mathrm{apex}}=-73.4^{\circ}$; see the all-sky projection in Fig. 4 for the direction of the apex on-sky (pink cross).
@@
We look to estimate the MW analogue mass, LMC analogue mass and the dynamical friction strength with knowledge of the LMC present-day position, velocity and the reflex motion parameters. Using our amortised posterior this can be efficiently evaluated for the given data. In Fig. 4 we show the joint, and individual, posterior distributions. Interestingly, our SBI framework is able to recover the true LMC mass from the cosmological MW–LMC analogue within the $16^{\rm th}$ and $84^{\rm th}$ percentiles. However, the mean of the LMC mass posterior distributions is biased towards higher values. Moreover, the dynamical friction strength is well constrained to values that are $\sim 2-3$ times greater than, and intriguingly not consistent with, those expected by the classical Chandrasekhar dynamical friction (i.e., $\log_{10}(\lambda_{\rm DF})=0$, Chandrasekhar1943). Similar to the deforming simulations, no specific value for the ‘true’ dynamical friction strength is quoted for this cosmological simulation, hence its absence from Fig. 4. Again we can use the orbit of this LMC analogue to find an approximate value given our treatment of dynamical friction which returns a value consistent with the posterior in Fig. 4. Exploring the ability to use our SBI framework to constrain dynamical friction remains of interest for future studies. Once more, the MW mass posterior is not well constrained away from the original prior distribution, implying that the supplied data points do not break model degeneracies.
@@ 5.1 Coverage probability test
Any posterior resulting from a generative model should be assessed for its accuracy through a variety of diagnostic tests before being used for inference given the actual observed data. The accuracy of an estimated posterior is often performed using some form of a coverage probability test. A coverage test in Bayesian analysis checks whether credible intervals have the expected probabilities (see, Jeffrey2025, sec. 2.4 for a concise explanation). In a 1-dimensional posterior setting, one can define a particular credible interval to be the narrowest interval containing, for example, $90\%$ of the probability weight. The Bayesian inference procedure takes in some observed data, $D_{\rm obs}$, and determines a posterior distribution, $p(\theta|D_{\rm obs})$, and hence a credible interval for $\theta$. For a coverage test, one uses a test parameter, $\theta_{\rm test}$, selected from the prior, $p(\theta)$, as the input to a simulation that produces the corresponding output data point, $D_{\rm test}$. From this, one can derive a posterior, $p(\theta|D_{\rm test})$, and therefore a credible interval. If the inference process has been correctly implemented then the true test parameter value, $\theta_{\rm test}$, will fall in this credible interval, in this example, $90\%$ of the time. Repeating this test for many sampled $\theta_{\rm test}$, and varying confidence intervals, one can gain confidence that the estimated posterior distributions are indeed correct.
@@
To perform a coverage test on our SBI posteriors, we will follow the ‘Tests of Accuracy with Random Points’ (TARP, Lemos2023, see their figs. 1&2 for futher intuition). For our application of SBI, this test is relatively straightforward as we have many pre-existing simulations with an amortised inference scheme, i.e., each data evaluation is computationally cheap. TARP coverage probabilities test the accuracy of estimated posteriors by only using samples made from the posterior. This technique is similar to simulation based calibration (Talts2018) but extends the idea to the full-dimensional posterior space instead of being solely 1-dimensional. We use the implementation of TARP in the sbi python package (tejero-cantero2020sbi) and show our results in Fig. 5.
@@
A valid SBI posterior will have a coverage probability that is similar to any given credibility interval one chooses. Fig. 5 demonstrates that the expected coverage does indeed match the credibility level. This validates our neural posterior estimation as being truly representative of the probability that each of our model parameters has some true value. This can be further quantified in two ways. Firstly, we can compute the area between the ideal TARP curve and our posterior TARP curves for credibility intervals greater than $0.5$; namely the Area To Curve (ATC) value. This number should be close to 0, a value $\gg 0$ indicates an estimated posterior that is too wide, conversely, a value $\ll 0$ indicates that the estimated posterior is too narrow. Secondly, we can calculate the p-value of a Kolmogorov-Smirnov test. The null hypothesis is that the ideal TARP curve and our posterior TARP curve are identical. If this p-value is less than $0.05$, then this null hypothesis is rejected. For our posterior distributions of MW mass, LMC mass, and dynamical friction strength when there is knowledge of the LMC present-day position, velocity, and reflex motion parameters (purple) we report an ATC value of $0.27$ and a Kolmogorov-Smirnov p-value of $1.0$. These values suggest that we are not drastically over-/under-fitting and are not required to reject the posterior.
@@
Figure 5: Tests of Accuracy with Random Points: This diagnostic check acts as a coverage probability test. The probability of true values in the appropriate credible intervals matches the expected coverage probability. A bootstrapped $2\sigma$ uncertainty is shown as the shaded region. This validates our SBI posterior estimation such that the posterior is representative of the probability that each of our model parameters has some true value.
@@ 5.2 Posterior Predictive Check
Although not explicitly a diagnostic test, we carry out a ‘Posterior Predictive Check’ (PPC) to act as a complementary diagnostic test. This test makes use of the fact that if the inference is correct, then any generated data, $D_{\rm pp}$, using simulation parameter values sampled from the posterior, $\theta_{\rm pp}$, should be similar to the observed data, $D_{\rm obs}$ (Lueckmann21a). A PPC can provide an intuition about any bias introduced in inference e.g., determining whether or not the generated data systematically differ from the observed data used during the inference.
We carry out a PPC for $1,024$ samples from our posterior distributions of the MW mass, LMC mass, and dynamical friction strength when there is knowledge of the LMC present-day position, velocity, and knowledge of the reflex motion parameters. We repeat this for the three cases we consider in this work: the validation on a rigid simulation in Sec. 4.1, the validation on a deforming simulation in Sec. 4.2 and the validation on a cosmological simulation in Sec. 4.3. To do this, we sample values for the MW mass, LMC mass and dynamical friction strength from the posterior evaluated using data in these three scenarios. We then re-run rigid MW–LMC simulations adopting these parameter values and output the reflex motion parameters to compare to the original input reflex motion data. In Fig. 6, we visualise the resulting generated data as coloured $1\,\sigma$ and $2\,\sigma$ contours and the original data (coloured crosses). The 1D panels show the filled $1\,\sigma$ region. Clearly, the input reflex motion data points used to evaluate the posteriors fall well within the $2\,\sigma$ confidence intervals across all fidelity MW–LMC simulations we consider. Furthermore, for the rigid and deforming simulations, the input data is contained within their $1\,\sigma$ confidence interval. This suggests that our return SBI posteriors are not subject to any biases and are truly representative of the input data.
@@
5.3 Train–Test split
We carry out a test–train split check to determine whether the posterior is performing accurately across on the full parameter ranges. To achieve this, we reserve $5,000$ of our rigid simulations to act as the test set which are not used to train the neural posterior estimator. The posterior training set consists of all other simulations. In Fig. 7, we compare the true LMC mass to its predicted median value, with $16^{\rm th}-84^{\rm th}$ percentile error bars, for a random sample of $100$ rigid simulations, taken from the larger test set of $5,000$ simulations. The true and predicted LMC masses agree within uncertainty except for the very largest LMC masses, where the prior begins to bias the predicted mass to lower values. Nonetheless, this high mass region is unlikely to contain the true present-day LMC mass, making this deviation inconsequential.
Figure 7: True vs. predicted LMC mass for a test set of simulations: we show the true LMC mass compared to its predicted median value, with $16^{\rm th}-84^{\rm th}$ percentile error bars, for a test set of $100$ of our rigid simulations which have not been used to train the neural posterior estimator. The training set used to learn the amortised posterior distribution consists of all other simulations. Towards the largest LMC masses, the neural posterior estimator biases the estimated LMC mass towards lower values, driven by the prior probability coverage; see right panel.
@@ 6.1.1 Improvements to the Milky Way – LMC models
Firstly, we have assumed that the MW is at equilibrium prior to the LMC’s infall. The extent to which the MW has any ‘smooth’ stellar halo is still of interest. The inner halo is made up of many substructures from accreted dwarf galaxies, globular clusters and stellar streams, some of which are well-mixed in phase-space, whereas others retain a degree of coherent structure at present day (e.g., Naidu2020; Malhan2022; Dillamore2025). Similarly, the outer halo contains a plethora or accreted substructures (e.g., the merger with the Sagittarius dSph galaxy, Helmi2020; Vasiliev2021; Hunt2025, and references therein). Accounting for this induced state of disequilibrium in the MW prior to the LMC’s infall is an oversight in our current modelling. One way to incorporate this is to include an overall net compression velocity to the MW stellar halo due to its contraction over time (Chandra2025; Bystrom2025). Additionally, we are not required to account for masking substructures that are seen in the observed MW stellar halo (e.g., the Sagittarius stellar stream, Amarante2024; Bystrom2025).
@@
Secondly, we assumed that the LMC has recently made its first pericentric passage. However, there is the alternative scenario where the LMC is on its second pericentric passage (Vasiliev2024). In this case, the first pericentric passage would cause a degree of disequilibrium in the MW, which we do not account for. However, the first pericentric passage would have occurred around $\sim 5-10\,\rm Gyr$ ago and at a distance $\geq 100\,\rm kpc$. Hence, almost all dynamical disequilibrium induced by any previous passage will have been superseded by its most recent one (Vasiliev2024; Sheng2024).
@@
Other modelling improvements include accounting for any mass changes of either the MW or LMC, triaxiality in the MW stellar haloes and/or distortions. The mass changes are especially important when considering cosmological analogues as both the host and satellite masses can change by a significant fraction throughout their interaction. In turn, this affects the degree of dynamical friction that the infalling satellite is subject to. In our current analytic dynamical friction prescription, this makes it difficult to separate out the effect of a time varying satellite mass and the inferred value of the dynamical friction strength at present-day, $\lambda_{\rm DF}$ (see Sec. 2.1.3 and Koposov2023). A further future model improvement would be to implement a more careful treatment of time dependent dynamical friction to account for this. Moreover, Chandrasekhar’s formula becomes increasingly ill-posed when considering two body systems of comparable masses (White1983; Binney2008; Jethwa2016). Finally, $v_{\rm travel}$ is sensitive to the mass enclosed by the LMC within a given distance rather than the total LMC mass which involves an extrapolation. We aim to reflect this in future work to return posteriors on the LMC mass enclosed.